OpenStudy (amurciano2):

Calculate the mass of CO2 emitted into the atmosphere upon accelerating the SUV from 0.0 mph to 64.0 mph . Assume that the required energy comes from the combustion of octane with 30% efficiency. (Hint: Use KE=1/2mv2 to calculate the kinetic energy required for the acceleration.) A large sport utility vehicle has a mass of 2400 kg . Express your answer using two significant figures.

OpenStudy (aaronq):

Youre going from 0 miles/h to 64 miles/h, this is the change in velocity \(\sf \Delta KE=\dfrac{1}{2}m(v_f^2-v_i^2)=\dfrac{1}{2}mv_f^2\) Before you plug in the values you need to do some conversions - i suggest you change miles/h to m/s, then your units work out to joules. After, you need take into account the efficiency of 30%. \(\sf \dfrac{KE_{needed}}{KE_{actual}}=\dfrac{0.3}{1}\rightarrow KE_{actual}=\dfrac{KE_{needed}}{0.3}\)